K is between J and M. L is between K and M. M is between K and N. If Jn=30, KM=8, and Jk=kL=lm, what is MJ?
1
Expert's answer
2020-09-15T16:17:19-0400



"KL+LM=KM"

If "KL=LM" and "KM=8", then

"2*KL=8"

"KL=8:2"

"KL=4"

If "KL=JK=LM", then "KL=JK=LM=4"

So, "MJ=JK+KL+LM"

"MJ=4+4+4"

"MJ=12"

And this task has surplus statement "JN=30", that was unnecessary to solve.

Solution: "MJ=12"